Find the value of K such that the equalities are equivalent
5/11 x+5=2/11 x-16
5/11 x=2/11 x+k



Answer by stanbon(75887)

Find the value of K such that the equalities are equivalent
5/11 x+5=2/11 x-16
5/11 x=2/11 x+k
------
(5/11)x = (2/11)x - 21
(5/11)x = (2/11)x + k
----
k = -21
-------------
Cheers,
Stan H.
------------